A jetliner, traveling northward, is landing with a speed of 62.9 m/s. Once the jet touches down, it has 736 m of runway in which to reduce its speed to 11.1 m/s. Compute the average acceleration (magnitude and direction) of the plane during landing (take the direction of the plane's motion as positive).
Answer #1

initial speed = V0 = 62.9 m/sec ( north)

final speed = V = 11.1 m/sec

displacement = S = 736 m

V^2 - V0^2 = 2 a s

acceleration = a = ( V^2 - V0^2) / 2S

                     a = 11.1^2 - 62.9^2 / 2 * 736

                     a = - 2.6 m/sec^2

magnitude = 2.6 m/sec^2

direction is towards south